OpenStudy (anonymous):

simplify: cos(6x)cos(2x)-sin(6x)sin(2x)

OpenStudy (anonymous):

Go to your compound angle formulae; this one will stand out to you blatantly.

myininaya (myininaya):

cos(6x+2x)

myininaya (myininaya):

cos(8x)

OpenStudy (anonymous):

how do you simplify this though? i am having trouble with the steps. and what is the compound formula? i don't believe we went over that in class.

myininaya (myininaya):

cos(8x) is simplified. cos(u-v)=cos(u)cos(v)+sin(u)sin(v) and cos(u+v)=cos(u)cos(v)-sin(u)sin(v) is a couple of the formulas

OpenStudy (anonymous):

okay, thank you so much!